OpenStudy (anonymous):

solve for x. 1/25=125^(3x-5). a. 17/9 b. 1/3 c. 13/9 d. 1 e. none of the above

OpenStudy (anonymous):

125=5^3 1/25=5^-2 therefore 3x-5 must equal -5 3x-5=-5 3x=0 x=0 therefore the answer is e.

jhonyy9 (jhonyy9):

-2=3(3x-5) -2=9x-15 13=9x x=13=9 so c. is right

jhonyy9 (jhonyy9):

sorry x=13/9